# Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data

### v2

@@ 5.3. Application: Search & Ranking
We then synthetically generated HellaSwag-bench by prompting Llama-70B to create 200 paper abstracts, each with a specified accuracy value, randomly sampled from . These datasets allows us to evaluate LOTUS’ LM-based sorting algorithms on a task with objective ground truth an ranking criteria other than relevance-based ones. We note that an alternative approach to these tasks could efficiently leverage sem_map to extract accuracy values on abstracts from either dataset, then perform a structured sort. However, our evaluation focuses on assessing the semantic ranking capabilities of various top-k algorithms, and we find this benchmark useful for understanding performance trade-offs, which may generalize to a wider set of reasoning-based ranking queries. We report results for trials at temperature , similar to prior works (Khattab et al., 2023)
| Method | nDCG@10 | ET (s) |
| Search | 0.712 | 0.009 |
| Search + Reranker | 0.741 | 2.64 |
| LOTUS Top-k - Llama-70B | 0.775 | 33.6 |
| LOTUS Top-k - GPT-4o | 0.800 | 11.2 |
| Method | CIFAR | HellaSwag | ||
| nDCG@10 | ET (s) | nDCG@10 | ET (s) | |
| Search | 0.252 | 0.008 | 0.119 | 0.008 |
| Search + Reranker | 0.001 | 2.57 | 0.461 | 2.36 |
| LOTUS Top-k - Llama 70B | 0.710 | 39.6 | 0.975 | 63.6 |
Baselines. We consider two simple baselines: a semantic search, using the E5Model (Wang et al., 2024) for retrieval, and a search with re-ranking, using the E5Model for retrieval and the MixedBread cross-encoder (noa, 2024) for re-ranking. We use choose these models because they are recent high quality embedding and re-rankers.
LOTUS Program. Our LOTUS program performs a semantic top- over the documents.
The langex for the semantic top- on SciFact sorts based on relevance to the given claim, while the langex for the CIFAR-bench and HellaSwag-bench datasets sort abstracts by accuracy performance on the respective datasets. For SciFact, we perform a semantic search using the E5Model as the retriever to obtain 100 articles, before ranking them with the sem_topk. We report results using both Llama-70B and GPT-4o as the LM.
| Method | Scifact | CIFAR | HellaSwag | ||||||
| nDCG@10 | ET (s) | # LM Calls | nDCG@10 | ET (s) | # LM Calls | nDCG@10 | ET (s) | # LM Calls | |
| Quadratic Sort | 0.836 | 712 | 4950 | 0.868 | 634 | 4950 | 0.966 | 1,803 | 19,900 |
| Heap Top-k | 0.776 | 65.0 | 216 | 0.832 | 99.6 | 350 | 0.907 | 98.9 | 415.2 |
| QuickSelect Top-k | 0.776 | 42.4 | 285 | 0.746 | 41.3 | 303.95 | 0.909 | 59.1 | 620.95 |
| QuickSelect Top-k + Semantic Index | 0.775 | 33.6 | 229 | 0.710 | 39.6 | 307.7 | 0.975 | 63.6 | 672.7 |
Results. Tables LABEL:tab:ranking_scifact and 5 demonstrate the effectiveness of LOTUS’ semantic top- for complex search tasks. In addition, Table 6 highlights the rich implementation design space that this task presents. We walk through several note-worthy findings. We first turn our attention to Table LABEL:tab:ranking_scifact, which presents the results for each bench-marked method on the SciFact dataset, which uses a relevance-based sorting criterion. As expected, both semantic search programs with and without re-ranking present strong baselines. Specifically, the re-ranker model, which is a supervised model trained specifically for the task of relevance-based ranking, increases nDCG@10 by 3 percentage points, while trading off query efficiency compared to the semantic search baseline. Notably, the unsupervised LM-based LOTUS programs outperform the supervised re-ranker’s result quality. The table shows LOTUS semantic top-k program with Llama-70B and GPT-4o, which outperform the semantic search baseline by 6 and 9 percentage points respectively. The LOTUS programs with Llama-70B and GPT-4o also outperform the re-ranker by 3 and 5 points respectively, improving upon the quality of the supervised baseline. As expected, the improved result quality comes at a trade-off to query efficiency due to the cost of LM calls. Notably, LOTUS’ versatility allows programmers to easily compose each of these query pipelines and trade-off result quality and efficiency depending on application-specific requirements.
Turning our attention to Table 5, we study LOTUS’ generality in supporting arbitrary language-based ranking criteria over the dataset. On the CIFAR-bench and HellaSwag-bench datasets, which use a complex sorting criteria, we see that both semantic search baselines with and without re-ranking provide poor result quality with consistently low nDCG@10. The LOTUS program, using a semantic top- with Llama 70B, acheives significant accuracy gains, with and higher nDCG@10 than the best performing baseline on CIFAR-bench and HellaSwag respectively. These accuracy gains reflect the powerful reasoning capabilities of LMs efficiently orchestrated over the data. As expected, these significant accuracy gains come at an increase to execution time.
@@
We now analyze the efficiency of our semantic top- implementation along with it’s proposed optimizations compared to alternatives. Table 6 compares several semantic top- algorithms, namely a quick-select top- algorithm, a quick-select top- that leverages the similarity index for pivot selection, a heap-based top- algorithm, and a quadratic sorting algorithm. First, we see that the quadratic algorithm, which performs an LM comparison between each pair of input documents, offers consistently high result quality across each dataset. However, this method is prohibitively expensive, requiring more LM calls and over higher execution time than the alternative implementations. The heap top- and quick-select top- methods offer comparable result quality, but with interesting trade-offs in query efficiency. Notably the quick-select top- method offers lower execution time than the heap-based sorting method across all datasets, despite requiring more LM calls in some cases. This is because the quick-select top- implementation allows for efficient batch-processing in each round of the algorithm, whereas the heap-based top- incurs sequential LM calls during heap updates. For this reason, our current implementation leverages the quick-select top- algorithm, although we envision future iterations may leverage multiple algorithms and allow the user to declaratively trade-off accuracy, query throughput, and cost.
In addition to providing an efficient top- algorithm, the table also demonstrates our use of the similarity index for optimizing top- query performance. The quick-select top- algorithm optimized with the semantic similarity index for pivot selection demonstrates lower execution time at no accuracy loss on SciFact, where the ranking criteria correlates likely with semantic similarity. On the other hand, for the CIFAR-bench and HellaSwag-bench datasets, where the ranking criteria does not correlate with semantic similarity, we see that the similarity index has no significant impact on the accuracy or efficiency top- performance.
5.4. Application: ArXiv Paper Analysis
Dataset.
Lastly, we evaluate LOTUS’ use in analyzing a large corpus of documents, specifically focusing on the performance of sem_group_by and its approximation algorithm with accuracy guarantees. The dataset consists of 647 recent ArXiv articles we scraped, from the database (cs.DB), information retrieval (cs.IR), cryptography and security (cs.CR), and robotics (cs.RO) domains. The task is to group the papers according to the topics they discuss.
Baselines. The sem_group_by consists of two sub-tasks: (a) discovering the groups and labels, and (b) classifying each document using the discovered labels. As a baseline for the latter, we consider a semantic similarity join between the labels and each abstract. We use the E5 embedding model for retrieval.
IRIS Programs.
We use the LOTUS sem_group_by operator to find 5 groups based on the topic abstracts. We use Llama-70B as the LM, the E5Model as the retrieval model, and 100 samples for the approximation algorithm described in Section 4.
| () Advancements in Recommender Systems and Multimodal Data Integration |
| () Advancements in Generative Information Retrieval Systems |
| () Advancements in Large Language Models for Various Applications |
| () Advancements in AI Security and Malware Detection Techniques |
| () Advancements in Robotic Navigation and Manipulation Techniques |
@@ 5.4. Application: ArXiv Paper Analysis
Results. We first qualitatively analyze the labels discovered, in the first phase of LOTUS’ implementation. As Figure LABEL:tab:groupby shows, the discovered labels intuitively align with the ArXiv domains included in the paper scrap, including important and timely topics related to recommendation systems, retrieval, LLM applications, AI security, and robotics. Performing this unsupervised group discovery took 44.03 seconds, representing a tractable and efficient implementation.
@@
Finally, turning our analysis to the classification subprocedure in our sem_group_by implementation, we evaluate the declarative algorithmic approximation. Figure 11 compares the execution time and classification accuracy with respect to the oracle implementation for the operator implemented with varying accuracy targets using . We see that the oracle implementation, at , offers the slowest execution time. On the other hand, the similarity-join baseline has faster execution time, but only accuracy with respect to the LM. We find our proxy-based optimization offers diverse performance trade-offs and allows the user to declaratively interpolate between these two extremes. We also note that the sampling procedure used to perform this optimization took less than 5 seconds, representing a small relatives cost.
@@ 6. Related Work
Specialized LLM-based Relational Extensions. Several prior works extend relational languages with a set of logically row-wise LM-based operations to serve specialized tasks.
Palimpzest (Liu et al., 2024b) presents a declarative approach to data cleaning and extract-transform-load (ETL) tasks. The authors propose to automatically optimize relational operators with LM calls, and implement two row-wise relational operators, a newly proposed convert operator, which can be transparently optimized to perform entity extraction using LLMs, and an AI-based filter operation, logically similar to LOTUS’ sem_filter. The system also implements several query optimizations, such as operator re-ordering, model selection, and code synthesis and proposes several others as future work.
SUQL (Liu et al., 2024c) presents a SQL extension to support conversational agents with knowledge grounding over structured and unstructured data. Specifically, the system extends SQL with two new logically row-wise operators, answer, which prompts an LM to answer the user question over each row, and summary, which prompts an LM to provide a summary to the user over each row. The system can provide automatic optimizations, such as predicate re-ordering using a lazy evaluation approach and proposes to use retrieval to optimize some answer queries.
ZenDB (Lin et al., 2024) and EVAPORATE (Arora et al., 2023) tackle the task of automatically ingesting and extracting semi-structured documents into structured tables that can be queried using standard relational operators and languages. ZenDB extracts structure using a semantic hierarchical tree index, which is integrated with a SQL query engine to support efficient query processing over the extracted attribute values. The systems implements several optimizations, including predicate reordering, push-down, and projection pull-up. Additionally, EVAPORATE performs efficient entity extraction from semi-structured data using LM-based code generation and weak supervision to ensemble candidate functions.
@@
In contrast to these works, semantic operators are a general-purpose query model designed to capture broad-ranging applications, including both logically row-wise ones and more complex ones, such as joins, aggregation, ranking and search functions, as an extension of the relational model. For example, users can use sem_map for row-wise entity extraction over unstructured text fields, but can also perform other operations over entire tables, such as sem_join.
We also propose multiple execution algorithms and novel optimizations for these bulk operators (Section 4).
Some of the optimizations in prior specialized systems, such as lazy evaluation, operator reordering, model selection, and code synthesis, are worthwhile future work for LOTUS’ optimizer.
LLM UDFs.
Recent research (Liu et al., 2024a) and some existing DBMS vendors, such as
BigQuery (ver, [n. d.]),
Databricks (dat, [n. d.]) and Redshift (noa, 2023), offer LLM user-defined functions (UDFs) in SQL. LLM UDFs are a lower-level programming interface that is limited to row-wise LLM execution over the data, equivalent to our sem_map semantic operator. In contrast, our semantic operator programming model is high-level and declarative, allowing the system to automatically select which an algorithm to use and manage calls to the LM to support complex operations, including aggregations, ranking, and joins.
Liu et al. (Liu et al., 2024a) study how to optimize LLM UDF calls, demonstrating performance gains with a de-duplication and prefix-sharing method that reorders rows and input fields to maximize key-value (KV) cache reuse during LLM calls. This is orthogonal to our optimizations but could likely accelerate batch LLM calls from LOTUS.
LM Programming Frameworks. LM programming frameworks such as LangChain (lan, [n. d.]), LlamaIndex (lla, [n. d.]), and DSPy (Khattab et al., 2023), have gained significant popularity. They provide abstractions for programming with LMs, including utilities for creating prompts and post-processing LM outputs, and modules for common calling patterns such as RAG, chat-bots, and agents. DSPy also optimizes a pipeline of LLM calls to maximize a target metric by tuning prompts and model weights. However, all these frameworks focus on expressing a pipeline of LLM calls that takes a single small input (or a stream of inputs for chat), as opposed to programs for processing large datasets. While some of them support batched calls with multi-threading, support for bulk-processing is sparsely supported and largely un-optimized. In contrast, semantic operators and LOTUS are designed for bulk semantic processing and introduce novel operators and optimizations for this task.
@@
ML-based Query Processing. Many prior works study the use of machine learning (ML) in databases, but do not focus on LLMs, which present unique usage patterns and optimization opportunities. MADLib (Hellerstein et al., 2012) extends SQL with abstractions for machine learning and descriptive statistics. Prior works such as NoScope (Kang et al., 2017), TASTI (Kang et al., [n. d.]), SUPG (Kang et al., 2020), BlazeIt (Kang et al., 2019) and probabilistic predicates (Lu et al., 2018) propose methods to optimize queries involving expensive ML models over large datasets, typically in video analytics. Some optimizations proposed in these works, such as model cascades and predicate re-ordering, are also useful for optimizing LOTUS pipelines with language models. However, our bulk semantic processing setting requires new operators with significantly different semantics, such as sem_join and sem_groupby, which were not optimized in this prior work. These motivate our novel optimizations that can approximate these operators with statistical guarantees (Section 4).
Text2SQL A large body of prior work looks at Text2SQL (Yu et al., 2018; Yaghmazadeh et al., 2017; Zhang et al., 2024; Zelle and Mooney, 1996), using LMs to translate natural language queries to executable SQL code. These methods serve semantic queries that are expressible in traditioanl relational operators. By contrast, semantic operators provide new AI-based transformations, such as aggregations for summarizing large document sets, or rankings, joins and filters according to natural language expressions, which are not expressible in SQL. We note an interesting line of future work may build a natural language interface on top of semantic operators.

### v3

@@ 5.3. Search & Ranking
Relevance-based ranking has been widely studied in the context of information retrieval. In addition, our conversations with LOTUS users reveal a common need for ranking based on complex natural language criteria (e.g., ranking customer reviews based on how frustrated they sound). In this application, we assess LOTUS’ ranking capabilities on both types of ranking tasks with two datasets: BEIR’s SciFact test set (Thakur et al., 2021), a widely used benchmark for retrieval and re-ranking, and a new dataset, HellaSwag-bench, which we generated to assess more complex ranking tasks. For both we report average nDCG@10, a standard ranking metric, and execution time (ET). The SciFact dataset consists of scientific claims and a corpus of articles, where the task is to rank articles by relevance to a given claim. We sample 300 claims for our evaluation. HellaSwag-bench consists of 200 synthetic paper abstracts888Each paper abstract in HellaSwag-bench is synthetically generated by prompting Llama-70b to write a research abstract that claims a specified accuracy value, which we randomly sample from ., each reporting an accuracy on the HellaSwag dataset (Zellers et al., 2019), and the task is to rank abstracts according to their reported accuracy. This dataset provides an objective ground truth, while focusing on a reasoning-based ranking criteria, rather than a relevance-based one999While the Hellaswag-bench ranking task might be efficiently solved with a sem_map, to extract accuracy values, followed by a structured sorting, our evaluation focuses on assessing semantic ranking algorithms. We find this benchmark useful for understanding performance trade-offs, which may generalize to a wider set of reasoning-based ranking queries.. We report results for trials at temperature , similar to prior works (Khattab et al., 2023).
Baselines. In addition to the standard search baseline, we add a re-ranking method using the MixedBread cross-encoder (noa, 2024), a high quality re-ranker that is common in relevance-based retrieval. The AI UDF baseline performs a pointwise-ranking by prompting the LLM to assign a relevance score to each row, similar to prior work (Zhuang et al., 2024). The DocETL baseline uses the system’s LLM reduce operation for ranking, as described by the pre-print (Shankar et al., 2024).
Our LOTUS program uses the sem_topk operator with a semantic index on the document corpus of both datasets.
For the reranker, AI UDF, DocETL, and LOTUS programs on SciFact, we perform search to retrieve 100 articles, before re-ranking them with each method.
| Method | SciFact | HellaSwag-bench | ||||
| nDCG@10 | ET (s) | # LM Calls | nDCG@10 | ET (s) | # LM Calls | |
| Search | 0.712 | 0.009 | 0 | 0.119 | 0.008 | 0 |
| Reranker | 0.741 | 2.64 | 0 | 0.461 | 2.36 | 0 |
| AI UDF | 0.457 | 9.83 | 100 | 0.091 | 19.7 | 200 |
| DocETL - Llama 70B optimizer (avg. of 3 sucesses)* | N/A | N/A | N/A | 0.246 | 14.1 | 3 |
| DocETL - GPT 4o optimizer* | N/A | N/A | N/A | N/A | N/A | N/A |
| LOTUS | 0.765 | 36.3 | 213.2 | 0.919 | 57.0 | 506.4 |
* DocETL is unable to acheive any successful runs using GPT-4o for its optimizer on SciFact and HellaSwag-bench, nor using Llama-70B for its optimizer on SciFact
| Method | SciFact | HellaSwag-bench | ||||
| nDCG@10 | ET (s) | # LM Calls | nDCG@10 | ET (s) | # LM Calls | |
| Quadratic Topk | 0.768 | 801.2 | 4950 | 0.886 | 2116.6 | 19,900 |
| Heap Topk | 0.765 | 60.0 | 192.6 | 0.896 | 59.4 | 241.5 |
| Quickselect Topk | 0.771 | 40.5 | 237.0 | 0.893 | 49.4 | 448.1 |
| LOTUS | 0.765 | 36.3 | 213.2 | 0.919 | 57.0 | 506.4 |
Results.
We evaluated the performance of the LOTUS program against baselines, in comparison to the gold algorithm, and to alternative high-quality ranking algorithms we considered as gold algorithms for our sem_topk.
First Table 6 demonstrates that LOTUS achieves higher accuracy than the next best-performing baselines on Scifact and HellaSwag-bench. Specifically, as expected, the re-ranker offers competitive accuracy compared to the LOTUS program on Scifact, which focuses on relevance-based retrieval, the supervised task the re-ranker model was trained on. However, on HellaSwag-bench, the LOTUS program significantly outperforms the re-ranker on this reasoning-based task. We also observe that the point-wise AI UDF method and the list-wise ranking, used by DocETL, offer substantially lower accuracy, unable to outperform the re-ranking baseline. This reflects the limitations of point-wise and list-wise ranking, which has been studied in prior work (Qin et al., 2024). Notably, the DocETL baseline altogether fails to produce any successful execution plans using the GPT-4o optimizer on Scifact and Hellaswag-bench and using the Llama-70B optimizer on Scifact.
@@ Results.
Turning our attention to Table 7, we compare the quick-select top-k gold algorithm to several alternative candidates. The quadratic top-k, heap top-k, and quick-select top-k offer comparable accuracy, making each a viable gold algorithm. However, the three candidate choices provide significant differences in efficiency. The quadratic algorithm, which performs an LLM comparison between each pair of input documents, requires more LM calls and over higher execution time than the heap top-k and quick-select top-k. In addition, the heap top- and quick-select top- methods offer interesting trade-offs in LLM call complexity and execution time. Notably the quick-select top- offers lower execution time than the heap-based sorting method, despite requiring more LLM calls in some cases. This is because the quick-select top- implementation allows for efficient batched processing in each round of the algorithm, whereas the heap-based top- incurs sequential LLM calls during heap updates. LOTUS currently implements the quick-select topk- by default, favoring lower execution time, although we future iterations may leverage multiple algorithms.
Finally, the Table 7 compares the quick-select top-k with the LOTUS quick-select top-k using embedding-based pivot selection. We demonstrate that the optimization is lossless, and can decrease latency by . Specifically, this optimization is useful where document rank correlates with semantic similarity, which is likely on the Scifact dataset.
| () | Advancements in Recommender Systems and Multimodal Data Integration |
| () | Advancements in Generative Information Retrieval Systems |
| () | Advancements in Large Language Models for Various Applications |
| () | Advancements in AI Security and Malware Detection Techniques |
| () | Advancements in Robotic Navigation and Manipulation Techniques |
5.4. Topic Analysis on ArXiv Papers
A common unsupervised discovery task requires grouping a large document corpus by topics and assigning descriptive labels to each group. We consider this task over a dataset of 647 recent ArXiv articles, which we scraped from the database (cs.DB), information retrieval (cs.IR), cryptography and security (cs.CR), and robotics (cs.RO) domains. We aim to discover key topic labels.
We can succinctly represent this task using the sem_group_by operator with LOTUS.
AI UDFs, DocETL and standalone search systems do not serve operations for unsupervised semantic-based group discovery. While UQE (Dai et al., 2024) studies a tractable optimization for serving non-LLM aggregations, e.g., COUNT, with LLM-based group-bys using stratified sampling, it does not focus on implementing, optimizing, and evaluating a standalone group-by operator, but suggests a standalone group-by implementation similar to our sem_group_by gold algorithm. Thus, in this section, we focus our analysis on understanding the performance of our sem_group_by gold algorithm and analyzing our optimization with respect to it. We use Llama-70B and E5 embeddings. For approximations, we use a sample size of 100 and .
Results.
The sem_group_by consists of two sub-tasks: (a) discovering representative group labels, and (b) classifying each document to a discovered label.
@@
We first qualitatively analyze the labels discovered, in the first sub-task of the LOTUS implementation, which is equivalent to gold algorithm. As Figure 7 shows, the discovered labels intuitively align with recent research topics from the cs.DB, cs.IR, cs.CR, and cs.RO ArXiv domains related to recommendation systems, retrieval, LLM applications, AI security, and robotics. Performing this unsupervised group discovery took 44.03 seconds, representing a tractable implementation.
@@
Turning to the classification sub-task, we compare the performance of the gold algorithm to our approximation, which uses an embedding-based proxy here. Figure 8 compares the execution time and classification accuracy of the oracle-based gold algorithm (green circle), our approximations using both the Llama-70B oracle LLM and embedding-based proxy at varied accuracy targets (blue stars), and the embedding-based proxy alone (red circle). We see that the proxy-only baseline is faster than the oracle, but about less accurate. Our approximation algorithm allows users to declaratively interpolate between these two extremes by varying the accuracy target. We also note that the sampling-based optimization procedure used to perform this optimization took less than 5 seconds, representing a small relative cost.
5.5. Accuracy Guarantees Evaluation
Our approximation methods are designed to adaptively learn when to leverage a proxy to provide an accuracy guarantee, despite variations in proxy accuracy. To demonstrate the accuracy guarantees provided by our semantic operator optimizations, we study the sem_filter performance on the fact-checking performance with varied proxy models, recall and precision targets and failure probability parameters. Figure 9(a) and Figure 9(b) demonstrate the average accuracy observed over 20 trials for varied recall targets and precision targets. We evaluate the filter approximation with both the Llama-8B proxy (solid lines) and a weaker proxy, with TinyLlama-1B (dashed lines) with a failure probability of (red) and (blue). As expected, the average recall and precision values increase with increasing accuracy targets for either metric. Figure 9(c) demonstrates that decreasing the recall and precision target tends reduce the number of calls routed to the oracle. As expected, at a fixed accuracy target, our approximation requires more oracle calls when using the weaker proxy, in this case TinyLlama. Finally, in Figure 9(d) we record the percentage of 50 trials that do not meet the recall and precision target, both set to 0.9, using the TinyLlama proxy. The results confirm that the observed failure probabilities increase with the user-configured delta values, and are lower than the configured failure probabilities, as expected due to our conservative implementation.
@@ 6. Related Work
Data Systems for Row-wise LLM Operations.
Many prior works focus on map-like, row-wise LLM-based operators, either with generic AI user-defined functions (UDFs) or specialized operators (Liu et al., 2024b; Lin et al., 2024; Arora et al., 2023; Liu et al., 2024c).
AI UDF systems (MotherDuck, [n.d.]; WilliamDAssafMSFT, 2024; noa, [n.d.]d; dat, [n.d.]; noa, 2023; Liu et al., 2024a) provide a low-level, non-declarative programming interface for row-wise LLM invocations. These systems also integrate vector search, equivalent to our sem_search or sem_sim_join, and support SQL queries that combine row-wise AI UDFs with non-LLM operations (e.g. average, count).
Alternatively, several recent works implement optimized row-wise LLM operations for specialized tasks, including data cleaning, extract-transform-load (ETL) and conversational agents.
ZenDB (Lin et al., 2024), EVAPORATE (Arora et al., 2023) and Palimpzest (Liu et al., 2024b) provide LLM-based functions for extracting semi-structured documents into structured tables, logically similar to our sem_extract, and Palimpzest also implements an LLM-powered filter, logically similar to our sem_filter.
SUQL (Liu et al., 2024c) augments SQL with a two logically row-wise operators, answer, for questions-answering over each row, and summary, for row-wise LLM-based summarization, to serve knowledge-grounded conversational agents.
@@
In contrast to these works, semantic operators are a general-purpose query model for semantic transformations over datasets, including both logically row-wise ones and more complex ones (e.g., joins, aggregations and ranking).
Our detailed evaluation (Section 5) demonstrates the performance limitations of row-wise LLM execution models.
Best-Effort LLM-Based Analytics Systems. Several recent systems (Anderson et al., 2024; Dai et al., 2024; Shankar et al., 2024) go beyond row-wise LLM operators. However, these systems lack a formalism for LLM-based operators, providing no performance guarantees for their execution. Following a preprint (Patel et al., 2024a) of this work, DocETL (Shankar et al., 2024) proposes to use LLM agents as the query optimizer over LLM-powered operators. This is a promising direction for future work; however, our evaluation (Section 5) demonstrates that DocETL’s agentic approach leads to high variance in performance, no accuracy guarantees, and requires multiple reruns to achieve comparable accuracy to ours, with approximately 1 in 3 runs failing due to optimizer errors, indicating that more work is needed to make these methods robust. UQE (Dai et al., 2024) also studies optimizations for LLM-powered operators, proposing an embedding-based approximation for LLM-powered filters. In contrast to our methods, which adaptively learn how to leverage a given proxy to meet accuracy targets, UQE’s approach provides best-effort performance with no accuracy guarantees. UQE also studies optimizations for non-LLM aggregations (e.g., average, count) combined with LLM-based filter and group-by operators using stratified sampling, which is complementary to this work.
@@
General ML-based Query Processing. Prior works study the use of non-LLM machine learning (ML) in databases. In contrast, our work integrates LLMs, which motivates our new formalisms for language-based operators and novel optimizations. MADLib (Hellerstein et al., 2012) extends SQL with abstractions for descriptive statistics and machine learning (e.g., regression and classification). NoScope (Kang et al., 2017), TASTI (Kang et al., [n.d.]), SUPG (Kang et al., 2020), BlazeIt (Kang et al., 2019) and probabilistic predicates (Lu et al., 2018) propose methods to optimize queries involving expensive ML-based predicates (e.g., using object detectors) over large datasets, typically in video analytics. Some optimizations proposed in these works, such as model cascades and predicate re-ordering, are also useful for optimizing LOTUS pipelines with language models.
Table Question Answering. A large body of work, including retrieval-augmented generation (Lewis et al., 2021) and Text2SQL (Yu et al., 2018; Yaghmazadeh et al., 2017; Zhang et al., 2024; Zelle and Mooney, 1996), serves natural language interfaces for question-answering over databases. Typically, the LLM system invokes external tools, e.g., search APIs or SQL programs. Interestingly, these agentic workflows can instead invoke semantic operator programs as the underlying API for data processing. Recent work (Biswal et al., 2024) demonstrates the promise of this approach to outperform baseline TableQA methods.